# Superquantile-Gibbs Relaxation for Minima-selection in Bilevel Optimization

### v2

@@ Appendix D Proofs of results in Section˜5
$\psi_{\rho}(\theta_{i+1})\leq$
$\psi_{\rho}(\theta_{i})-\left[\tilde{\eta}_{i}-\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\right]\cdot\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
@@
$\leq\psi_{\rho}(\theta_{i})-\left[\tilde{\eta}_{i}-\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\right]\cdot\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$\quad+\tilde{\eta}_{i}\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2},$ (66)
@@
$\psi_{\rho}(\theta_{i+1})\leq$
$\psi_{\rho}(\theta_{i})-\left[\tilde{\eta}_{i}-\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\right]\cdot\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$\quad-\tilde{\eta}_{i}\langle\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+\tilde{\eta}_{i}\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2},$ (67)
@@
$\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\leq\frac{2[\psi_{\rho}(\theta_{i})-\psi_{\rho}(\theta_{i+1})]}{\tilde{\eta}_{i}}-2\langle\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$\quad-2\langle\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+2\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}.$ (68)
@@
Then by using Young’s inequality ($\langle{\bf a},{\bf b}\rangle\leq c\|{\bf a}\|^{2}/2+\|{\bf b}\|^{2}/(2c)$ for the vectors ${\bf a},{\bf b}$ and positive constant $c$), we have
$\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\leq\frac{2[\psi_{\rho}(\theta_{i})-\psi_{\rho}(\theta_{i+1})]}{\tilde{\eta}_{i}}-2\langle\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$\quad+4\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}+4^{-1}\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}+2\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}.$ (69)
@@
Using [li2022nonsmooth, Lemma D1], $\mathbb{E}_{u_{t}}[\mathsf{g}_{\rho}(\theta_{i})\mid\theta_{i}]=\nabla\psi_{\rho}(\theta_{i})$. Taking expectation given $\theta_{i}$
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}]\leq\frac{2[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+4\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+$
@@
$\quad+4^{-1}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}]+2\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\leq\frac{2[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+6\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+$
$\qquad\qquad\qquad\qquad\qquad+4^{-1}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}],$ (70)
where the last inequality comes from Jensen’s inequality and $\mathbb{E}_{u_{t}}[\mathsf{g}_{\rho}(\theta_{i})\mid\theta_{i}]=\nabla\psi_{\rho}(\theta_{i})$; $\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]=\mathbb{E}[\|\mathbb{E}_{u_{t}}[\mathsf{g}_{\rho}(\theta_{i})\mid\theta_{i}]-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]\leq\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$.
@@
$\overset{(b)}{\leq}2\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]+2\mathbb{E}[\|\tilde{\mathsf{g}}_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}|\theta_{i}].$ (71)
where (a) follows by $\|\mathcal{G}_{\Theta}(\theta,v_{1};\eta)-\mathcal{G}_{\Theta}(\theta,v_{2};\eta)\|\leq\|v_{1}-v_{2}\|$ and (b) is followed similarly to the last inequality in (70). Incorporating (70) in (71), we get
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]\leq\frac{4[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+14\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+$
$+2^{-1}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}].$ (72)
@@
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]\leq\frac{8[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+28\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}].$ (73)
@@
Step 2: Controlling the approximation error via Section˜5.1. We now invoke Section˜5.1 to bound the last term in (73):
@@
$\leq\frac{m^{2}}{\rho^{2}}\cdot\mathbb{E}[|\tilde{\psi}(\theta_{i}+\rho u_{t})-F_{\max}(\theta_{i}+\rho u_{t})|^{2}+|\tilde{\psi}(\theta_{i}-\rho u_{t})-F_{\max}(\theta_{i}-\rho u_{t})|^{2}]$
@@
$L=\mathcal{O}\left(\frac{\texttt{k}\mathrm{B}^{2}}{\epsilon_{v}^{2+2\texttt{k}}}\cdot\log(\epsilon_{v}^{-1})\right).$
By choosing $\tilde{\eta}_{i}=\eta\leq 1/L_{\psi,2}=m^{-1/2}\rho L_{F_{\max},1}^{-1}$ and integrating (74) into (73), followed by averaging over $\theta_{i}$, we obtain
$\frac{1}{N}\sum_{t=1}^{N}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho};\eta)\|^{2}]\leq\frac{4\psi_{\rho}(\theta_{0})-4\mathbb{E}[\psi_{\rho}(\theta_{N})]}{N\eta}+\frac{8m^{2}}{\rho^{2}}\cdot\mathcal{O}\left(\epsilon_{v}^{2}\right)$
$\leq\mathcal{O}\left(\frac{m^{\frac{1}{2}}\cdot(\psi_{\rho}(\theta_{0})-\mathbb{E}[\psi_{\rho}(\theta_{N})])}{N\rho}+\frac{m^{2}}{\rho^{2}}\cdot\epsilon_{v}^{2}\right)$
$\leq\mathcal{O}\left(\frac{m\cdot\text{diam}(\Theta)}{N\rho^{2}}+\frac{m^{2}}{\rho^{2}}\cdot\epsilon_{v}^{2}\right),$ (75)
@@
Let $\hat{\theta}:=\operatorname*{arg\,min}_{i\in[1:N]}\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho};\eta)\|$. Let us define $\hat{\theta}=\operatorname*{arg\,min}_{i\in[1:N]}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho};\eta)\|^{2}]$. Then by Jensen’s inequality, we have
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho};\eta)\|]=\mathcal{O}\left(\frac{m^{\frac{1}{2}}(\text{diam}(\Theta))^{\frac{1}{2}}}{\sqrt{N}\rho}+\frac{m}{\rho}\cdot\epsilon_{v}\right).$ (76)
Let us pick $\epsilon_{v}=\rho\epsilon m^{-1}$, and $N\geq m\cdot\text{diam}(\Theta)\cdot\epsilon^{-2}\rho^{-2}$. Then $\mathbb{E}[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho};\eta)\|]=\mathcal{O}\left(\epsilon\right)$. We also pick $b_{u}=\mathcal{O}(1)$. Then the total number of samples from Gibbs measure is
$N\times(b_{u}L+b_{u}|I|)=\mathcal{O}\left(\frac{m^{3+2\texttt{k}}\cdot\text{diam}(\Theta)\cdot(\texttt{k}\mathrm{B}^{2}+\tilde{C}_{f})\cdot\log\left(m\epsilon^{-1}\rho^{-1}\right)}{\epsilon^{2\texttt{k}+4}\rho^{2\texttt{k}+4}}\right).$ (77)

### v4

@@ Appendix D Proofs of Results in Section˜5
$\psi_{\rho}(\theta_{i+1})\leq\psi_{\rho}(\theta_{i})-\left[\tilde{\eta}_{i}-\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\right]\cdot\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}$
$-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
@@
$\leq\psi_{\rho}(\theta_{i})-\left[\tilde{\eta}_{i}-\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\right]\cdot\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}$
$-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+\tilde{\eta}_{i}\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2},$ (66)
@@
$\psi_{\rho}(\theta_{i+1})\leq\psi_{\rho}(\theta_{i})-\left[\tilde{\eta}_{i}-\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\right]\cdot\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}$
$-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$-\tilde{\eta}_{i}\langle\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+\tilde{\eta}_{i}\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2},$ (67)
@@
$\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\leq\frac{2[\psi_{\rho}(\theta_{i})-\psi_{\rho}(\theta_{i+1})]}{\tilde{\eta}_{i}}$
$-2\langle\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$-2\langle\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+2\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}.$ (68)
@@
Then by using Young’s inequality ($\langle{\bf a},{\bf b}\rangle\leq c\|{\bf a}\|^{2}/2+\|{\bf b}\|^{2}/(2c)$ for the vectors ${\bf a},{\bf b}$ and positive constant $c$, we have
$\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\leq\frac{2[\psi_{\rho}(\theta_{i})-\psi_{\rho}(\theta_{i+1})]}{\tilde{\eta}_{i}}$
$-2\langle\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+4\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}$
$+4^{-1}\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}+2\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}.$ (69)
@@
Using [54, Lemma D1], $\mathbb{E}_{u_{t}}[\mathsf{g}_{\rho}(\theta_{i})\mid\theta_{i}]=\nabla\psi_{\rho}(\theta_{i})$. Taking expectation given $\theta_{i}$
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}]\leq$
$\frac{2[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+4\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+$
$\quad+4^{-1}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}]+2\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}].$ (70)
@@
The term involving $\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})$ vanishes because $\mathbb{E}_{u_{t}}[\mathsf{g}_{\rho}(\theta_{i})\mid\theta_{i}]=\nabla\psi_{\rho}(\theta_{i})$ and $\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})$ is deterministic given $\theta_{i}$. Moreover, by $(a+b)^{2}\leq 2(a^{2}+b^{2})$,
$\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]\leq 2\,\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+2\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}].$
Substituting this into (70), we obtain
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}]\leq$
$\frac{2[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+4\,\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\qquad+8\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+4^{-1}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}].$ (71)
@@
$\leq 2\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]$
$\qquad+4\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}|\theta_{i}]+4\mathbb{E}[\|\tilde{\mathsf{g}}_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}|\theta_{i}].$ (72)
where (a) follows by $\|\mathcal{G}_{\Theta}(\theta,v_{1};\eta)-\mathcal{G}_{\Theta}(\theta,v_{2};\eta)\|\leq\|v_{1}-v_{2}\|$. Incorporating (71) in (72), we get
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]\leq\frac{4[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}$
$+12\,\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+20\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\qquad+2^{-1}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}].$ (73)
@@
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]\leq$
$\frac{8[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+24\,\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\qquad+40\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}].$ (74)
@@
Step 2: Controlling the directional and oracle errors. We first bound the variance coming from the random directions. Since the variables $\{\mathsf{g}_{\rho}(\theta_{i},u_{i,t})\}_{t=1}^{b_{u}}$ are i.i.d. and satisfy $\mathbb{E}_{u_{t}}[\mathsf{g}_{\rho}(\theta_{i},u_{i,t})\mid\theta_{i}]=\nabla\psi_{\rho}(\theta_{i})$, we have
$\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$=\frac{1}{b_{u}^{2}}\sum_{t=1}^{b_{u}}\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i},u_{i,t})-\nabla\psi_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\leq\frac{1}{b_{u}}\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i},u)\|^{2}\mid\theta_{i}]$
$\leq\frac{m^{2}L_{F_{\max},1}^{2}}{b_{u}},$ (75)
where we used
$\|\mathsf{g}_{\rho}(\theta_{i},u)\|=\frac{m}{2\rho}\,|F_{\max}(\theta_{i}+\rho u)-F_{\max}(\theta_{i}-\rho u)|\leq mL_{F_{\max},1}.$
We now invoke Section˜5.1 to bound the oracle term in (D.3):
@@
$\leq\frac{m^{2}}{\rho^{2}b_{u}}\sum_{t=1}^{b_{u}}\mathbb{E}[|\tilde{\psi}(\theta_{i}+\rho u_{i,t})-F_{\max}(\theta_{i}+\rho u_{i,t})|^{2}$
$\hskip 91.04872pt+|\tilde{\psi}(\theta_{i}-\rho u_{i,t})-F_{\max}(\theta_{i}-\rho u_{i,t})|^{2}]$
@@
$L=\mathcal{O}\left(\frac{\texttt{k}^{2}\mathrm{B}^{2}}{\epsilon_{v}^{2+2\texttt{k}}}\cdot\log^{2}(\epsilon_{v}^{-1})\right).$
By choosing $\tilde{\eta}_{i}=\eta\leq 1/L_{\psi,2}=m^{-1/2}\rho L_{F_{\max},1}^{-1}$ and integrating (D.3) and (76) into (D.3), followed by averaging over $\theta_{i}$, we obtain
$\frac{1}{N}\sum_{t=1}^{N}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho};\eta)\|^{2}]\leq\frac{8\psi_{\rho}(\theta_{0})-8\mathbb{E}[\psi_{\rho}(\theta_{N})]}{N\eta}+\mathcal{O}\left(\frac{m^{2}}{b_{u}}\right)$
$\qquad\qquad+\frac{40m^{2}}{\rho^{2}}\cdot\mathcal{O}\left(\epsilon_{v}^{2}\right)$
$\leq\mathcal{O}\left(\frac{m^{\frac{1}{2}}\cdot(\psi_{\rho}(\theta_{0})-\mathbb{E}[\psi_{\rho}(\theta_{N})])}{N\rho}+\frac{m^{2}}{b_{u}}+\frac{m^{2}}{\rho^{2}}\cdot\epsilon_{v}^{2}\right)$
$\leq\mathcal{O}\left(\frac{m\cdot\text{diam}(\Theta)}{N\rho^{2}}+\frac{m^{2}}{b_{u}}+\frac{m^{2}}{\rho^{2}}\cdot\epsilon_{v}^{2}\right),$ (77)
@@
Let us define $\hat{\theta}=\operatorname*{arg\,min}_{i\in[1:N]}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho};\eta)\|^{2}]$. Then by Jensen’s inequality, we have
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho};\eta)\|]=\mathcal{O}\left(\frac{m^{\frac{1}{2}}(\text{diam}(\Theta))^{\frac{1}{2}}}{\sqrt{N}\rho}+\frac{m}{\sqrt{b_{u}}}+\frac{m}{\rho}\cdot\epsilon_{v}\right).$ (78)
Let us pick $\epsilon_{v}=\rho\epsilon m^{-1}$, $N\geq m\cdot\text{diam}(\Theta)\cdot\epsilon^{-2}\rho^{-2}$, and $b_{u}\geq m^{2}\epsilon^{-2}$. Then $\mathbb{E}[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho};\eta)\|]=\mathcal{O}\left(\epsilon\right)$. We also pick $b_{u}=\mathcal{O}(m^{2}\epsilon^{-2})$. Then the total number of samples from Gibbs measure is
$N\times(b_{u}L+b_{u}|I|)=\mathcal{O}\left(\frac{m^{5+2\texttt{k}}\cdot\text{diam}(\Theta)\cdot(\texttt{k}^{2}\mathrm{B}^{2}+\tilde{C}_{f})\cdot\log^{2}\left(m\epsilon^{-1}\rho^{-1}\right)}{\epsilon^{2\texttt{k}+6}\rho^{2\texttt{k}+4}}\right).$ (79)